I don't know how to set this up:
Bill leaves his house for Maya's house riding his bicycle at 8 miles per hour. At the same time, Maya leaves her house heading toward Bill's house walking at 3 miles per hour. Write a system of equations to represent the distance, d, each is from Maya's house in h hours. They live 8.25 miles apart.

Answer by edjones(8007)

st=d s=speed, t=time, d=distance
How long did it take until they met?
d=distance from Maya's house.
t=d/3 for Maya
t=(8.25-d)/8 for Bill
d/3 = (8.25-d)/8 Since the times are equal.
8d=24.75-3d Multiply each side by 24 to eliminate fractions.
11d=24.75
d=2.25 mi from Maya's house.
.
Ed